# Current distribution and chirality in the regime of the quantized Hall effect

### v1

@@ II.2 Coherent and chiral model of $n$-terminal device
Next, we derive a general model for the current in a 2DES device prepared in the coherent and chiral limit but equipped with $n$ ohmic contacts, cf. Fig. 3. The model intends to describe current flow through a Hall bar in the region of a quantized plateau. The inner circle in Fig. 3 indicates the coherent 2D conductor, the radial lines symbolize the $n$ terminals $i=1,2,...,n$. External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the coherent part of the sample and each contact. The current $I_{i}$ flows through the $i$th contact with ohmic resistance $R_{i}$ to ground.
To describe the coherent inner part of the sample, we employ the Landauer-Büttiker formalism and write
@@
where $G_{ij}$ are the coherent conductance coefficients, which can be formally derived from the scattering matrix of the problem. Assuming chiral transport, current flows only between adjacent contacts. The direction of the chirality is linked to the direction of the magnetic field vector. In our case as sketched in Fig. 3, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}=1$ and $G_{ji}=0$ for $i=j+1$. The unitarity of the scattering matrix implies $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our chiral conductivity tensor takes the form
@@
where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels. The coherent limit applies for the quantized plateaus of the Hall resistance, where $\nu$ is an integer. Inserting Eq. (6) into Eq. (5) and solving the equation system yields the currents
$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)$ (7)
where we introduced
@@
Equation (7) describes the current distribution into the $n$ ohmic contacts of a chiral and coherent sample. For $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$ the currents can be written as
$I_{i>1}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,I_{1}\,,$ (10)
which are independent of $R_{1}$ as expected.
II.3 Drude model of $n$-terminal device
Our model describing coherent currents is based on the unitarity of the scattering matrix. However, regarding the actual macroscopic currents into various contacts, the unitarity condition implies the same result as we find from applying Kirchhoff’s laws. Therefore, if we replace $R_{\text{H}}$ by $(\nu G_{0})^{-1}$, for $n=3$ and $V_{2}=V_{3}=0$, Eq. (7) of the coherent and chiral model predicts the identical currents $I_{2,3}$ as the transversal current component $I_{xy}^{(2,3)}$ of the Drude model discussed in Sec. II.1 above. By extrapolating this observation, we can expand our prediction of the Drude model for the current distribution of an $n$-terminal junction and find
$I_{i}$ $=$ $\frac{1}{1+(\mu B)^{2}}I_{xx}^{(i)}+\frac{(\mu B)^{2}}{1+(\mu B)^{2}}I_{xy}^{(i)}\quad\text{with}$
$I_{xx}^{(i)}$ $=$ $\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (11)
$I_{xy}^{(i)}$ $=$ $\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,.$
The longitudinal component, $I_{xx}^{(i)}$, is determined by Ohm’s law while the transversal component, $I_{xy}^{(i)}$, is identical to Eq. (7), where we use $R_{\text{H}}=(\nu G_{0})^{-1}$.
To correctly describe the regions of the plateaus of a quantized Hall resistance, we use constant and integer $\nu$ and assume $\mu=\infty$, a property of ballistic transport, such that $I_{xx}^{(i)}=0$. Then Eq. (II.3) is identical to Eq. (7).
@@ II.3 Drude model of $n$-terminal device
Multiterminal current measurements yield the distribution of the current into the different contacts. In this section, we compare the results of multi-terminal measurements and theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$. We will see, that multiterminal measurements can provide answers to the questions about chirality and the local current distribution in magnetotransport.
@@ III.1 Sample and measurement setup
Figure 5: Three-terminal measurement with symmetric circuit, cf. inset in (a). A voltage of $V=1\,$mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating. $\vec{B}$ points from the surface into the 2DES plain such that $I_{xy}$ is bent towards contacts 2 and 3. (a) Current ratios $-I_{3}/I_{1}$ and $-I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq. (II.3) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq. (II.3) with $\mu(B)$ as free parameter (dotted red line). (b) enlargements near $\nu=1$ showing experimental data and the Drude model prediction for $\mu=\mu_{0}$.
Figure 6: Mobility $\mu(B)$ as determined from SdH-oscillations of the longitudinal resistance between contacts 2 and 3 (black dots) in comparison with $\mu(B)$ found by fitting Eq. (II.3) to the three-terminal measurements shown in Fig. 5 (red squares).
Figure 7: Four-terminal measurements for three different contact configurations, cf. insets, further configurations and labeling as in Fig. 5. (a) symmetric four-terminal circuit. (b) and (c) variations of asymmetric four-terminal circuits.
@@ III.2 Three-terminal measurements
First, we discuss a three-terminal measurement in a mirror symmetric geometry, for which we keep the contacts 2, 4 and 6 electrically floating; current flows from contact 1 into contacts 3 and 5. In Fig. 5(a), we present as solid (blue) lines the measured current ratios $-I_{3}(B)/I_{1}(B)$ and $-I_{5}(B)/I_{1}(B)$. As expected, we find that $|I_{3}|$ increases with $B$ while $|I_{5}|$ decreases accordingly, such that $I_{3}(B)+I_{5}(B)=-I_{1}(B)$. The dashed black line indicates the prediction of the Drude model according to Eq. (II.3) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$. For this large mobility, the Ohmic term (first term) of Eq. (II.3) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$). In this limit, with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$, Eq. (II.3) becomes identical to the coherent chiral model expressed in Eq. (7).
@@
The Drude model and the measured data agree in the classical limit for $B\lesssim 0.5$ T. As we increase $B$ further, the Drude model with a constant $\mu(B)=\mu_{0}$ still predicts the correct trend, but we observe growing deviations for filling factors away from integer values [top axis of Fig. 5(a)]. The enlargement around the first plateau of the QHE in Fig. 5(b) clarifies, that model and experiment match exactly only at specific values of $B$ including integer bulk filling factors (here $\nu=1$). An exact match between measurements and model can be achieved by using $\mu(B)$ as a free fit parameter in Eq. (II.3). For the dotted red line in Fig. 5(a), we used the magnetic field dependent values of $\mu(B)$ as shown by red squares in Fig. 6. For comparison, we have also determined $\mu(B)$ from the Shubnikov-de-Hass oscillations of a standard longitudinal resistance measurement, see the black dots in Fig. 6.  [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to $T>4\,$K between the two measurements. The agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq. (II.3) to our three-terminal current measurements supports the validity of our model.
@@
The two regions of divergent $\mu(B)$ in Fig. 6 correspond to the plateau regions at $\nu=2$ and $\nu=1$. Here, $I_{xx}=0$ and the coherent and chiral model expressed in Eq. (7) is valid. Away from the plateau regions, we find strongly reduced mobility values with $201\lesssim\mu<$ 20, which indicate diffusive transport and a finite ohmic contribution $I_{xx}$ to the current.
The overall agreement between the prediction of Eq. (II.3) and our three-terminal measurements is convincing. On the one hand, it supports the assumption of coherent and chiral transport for the quantized plateaus of the Hall resistance. On the other hand, it points to non-chiral transport described by the Drude model not only in the classical regime of the Hall effect but also for large $B$ in between the plateaus.
@@ III.3 Four-terminal measurements
In Fig. 7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements. Again, we measure the currents into contacts $i=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig. 7(a), contact 6 in (b), and contact 2 in (c). If we assume $\mu(B)=\mu_{0}$, Eq. (II.3) of the Drude model reproduces the overall dependences $I_{i}(B)$ with a similar accuracy as it did for the three-terminal measurement (dashed black lines). The dotted red lines correspond to the predictions of Eq. (II.3) using the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf. Fig. 6. Again, for $B\lesssim 0.5\,$T, theory and experiment agree well. For $B>0.5\,$T, the agreement is still perfect for the plateau regions but becomes worse away from the plateaus of quantized Hall resistance. experimental results is even worse (for $B>0.5\,$T).
This observation allows two immediate conclusions: (1) for the plateau regions, the chiral transport already found in the three-terminal measurement is confirmed; (2) away from the plateaus, the non-chiral transport is also confirmed. However, for more than three terminals, the model prediction of Eq. (II.3) fails for $B>0.5\,$T.

### v2

@@ II.1 Coherent and chiral model of $n$-terminal device
The inner circle in Fig. 2 indicates the 2DES, the radial lines symbolize the $n$ terminals $i=1,2,...,n$ connected via ohmic contacts to the 2DES. External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the 2DES and each contact. The current $I_{i}$ flows through the $i$th contact via the ohmic resistance $R_{i}$ to ground.
To describe the current flow through the 2DES (inner part of the sample), we employ the Landauer-Büttiker formalism (not to be confused with the LBP) and write
@@
where $G_{ij}$ are the conductance coefficients between mutual contacts. They can be formally derived from the scattering matrix of the problem. The left hand side of Eq. (2) is a general formulation of the transmission properties of the 2DES based on the Landauer-Büttiker formalism, while the last term of Eq. (2) applies Ohm’s law to the macroscopic contacts. If the transport within the 2DES is chiral, only those $G_{i,j\neq i}$ connecting adjacent contacts can be non-zero. The direction of the chirality is linked to the direction of the magnetic field vector. In our case sketched in Fig. 2, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}\neq 0$ for $i=j+1$. Assuming coherent transport, the scattering matrix is unitary, implying $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our coherent and chiral conductivity tensor takes the form
@@
where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels. While the assumption of coherence guaranties a unitary scattering matrix, for the validity of Eq. (3) it is actually sufficient to require the absence of momentum scattering. For the quantized plateaus of the Hall resistance $\nu$ is an integer. Inserting Eq. (3) into Eq. (2) yields a solvable linear equation system providing us with the prediction of the current into the $i$th contact
$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)\,,$ (4)
@@
A simplifying special case of some relevance is that of $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$, which yields the current ratios
$\frac{I_{i>1}}{I_{1}}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,.$ (7)
II.2 Drude model of an $n$-terminal device
The Drude model provides a solution of Eq. (2), which describes the classical limit of diffusive and incoherent currents inside the 2DES. Using $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ we can rewrite Eq. (1) in terms of the magnetic-field-dependent conductivity tensor as
where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$. For a laterally confined 2DES, in the steady state we expect $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
The structure of the conductivity tensor allows us to partition the non-vanishing current density $j_{x}$ into two components with distinct physical meanings, namely
$j_{x}\equiv j_{xx}+j_{xy}\quad\text{with}$
$j_{xx}=\frac{\sigma_{0}}{1+(\mu B)^{2}}\,E_{x}\quad\text{and}\quad j_{xy}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{\text{H}}}\,E_{y}\,.$
The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is the ohmic term, proportional to the specific conductivity and driven by the longitudinal electric field $E_{x}$. Caused by the additional Hall resistance at finite $B$, the ohmic current density is reduced by the factor $\left[1+(\mu B)^{2}\right]^{-1}$. The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is driven by the perpendicular electric field and attenuated by the non-local Hall resistance. For $\mu B\gg 1$ the ohmic component $j_{xx}$ vanishes and the current density $j\to j_{xy}$ becomes independent of $\sigma_{0}$. It implies, that momentum scattering is absent. At the same time, the transversal $j_{xy}$ is bent into a current carrying contact as it follows the charged edge of the Hall bar.
Figure 3: Drude model: Simplifying sketch showing the partitioning of the current contributions $I_{xx}$ versus $I_{xy}$ at a three-terminal junction. A negative voltage is applied to the left contact, while the other contacts are grounded. $I_{xx}$ is devided according to Kirchhoff’s current rule. Neglecting the partial reflection of current at ohmic contacts, only one egde is negatively charged, other edges are at ground potential. $I_{xy}$ is guided by the charged edge and flows entirely from the left to the lower contact.
In our experiments, we measure the current, which corresponds to the integral of the current densities in Eq. (II.2) across the Hall bar. For the partition of the current at a junction we can now use the following rules: The ohmic $I_{xx}$ branches according to Kirchhoffs current law. For simplicity, we neglect the small resistance $R_{0}$ of the 2DES ($R_{i}/R_{0}\sim 50$). The non-ohmic $I_{xy}$ partitions according to Eq. (4), if we include non-integer filling factors [$R_{\text{H}}=(\nu G_{0})^{-1}$]. Therefore, in the limit $\mu B\gg 1$, the prediction for the current division of the Drude model becomes identical to that of our coherent and chiral model derived in Sec. II.1. To illustrate the different behaviors of $I_{xx}$ and $I_{xy}$, in Fig. 3 we sketch the current division at a three-terminal junction. Combining both contributions, the Drude model predicts
$I_{i}$ $=$ $I_{xx}^{(i)}+I_{xy}^{(i)}\quad\text{with}$
$I_{xx}^{(i)}$ $=$ $\frac{1}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (10)
$I_{xy}^{(i)}$ $=$ $\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,$
for the current flowing into the $i$th contact.
The original scope of the classical Drude model was a description of diffusive transport at $\mu B\lesssim 1$. Interestingly, in the limit $\mu B\to\infty$, the first term of Eq. (II.2) vanishes, $I_{xx}\to 0$, and the prediction of the Drude model becomes identical to Eq. (4). This property also means that for the limits $\mu B\lesssim 1$ versus $\mu B\gg 1$, Eq. (II.2) precisely describes both of our limiting models. For intermediate $\mu B$, Eq. (II.2) smoothly interpolates between the limits. This is useful, because $\mu(B)$ strongly oscillates due to the Shubnikov-de-Haas oscillations.
@@ II.2 Drude model of an $n$-terminal device
Multiterminal current measurements yield the distribution of current into various contacts. In this section, we discuss the results of multi-terminal measurements and compare them with theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$. We will see, that even though multiterminal measurements do not provide the local current distribution across the Hall bar, their comparison with reasonable limiting models can provide answers to the questions about chirality of the current for the QHE.
@@ III.1 Sample and measurement setup
Figure 5: Three-terminal measurement with symmetric circuit, cf. inset in (a). A voltage of $V=1\,$mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating. $\vec{B}$ points from the surface into the 2DES plain such that $I_{xy}$ is bent towards contacts 2 and 3. (a) Current ratios $-I_{3}/I_{1}$ and $-I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq. (II.2) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq. (II.2) with $\mu(B)$ as free parameter (dotted red lines). (b) enlargements near $\nu=1$ showing experimental data, the Drude model prediction for $\mu=\mu_{0}$ and the model prediction of Eq. (4) for coherent and chiral current at $\nu=1$ (dotted red lines).
Figure 6: Mobility $\mu(B)$ as determined from SdH-oscillations of the longitudinal resistance between contacts 2 and 3 (black dots) in comparison with $\mu(B)$ found by fitting Eq. (II.2) to the three-terminal measurements shown in Fig. 5 (red squares).
Figure 7: Four-terminal measurements for three different contact configurations; the blue numbers in the insets indicate grounded contacts, black numbers indicate floating contacts and the external voltage is applied to contact 1 (red); further details as in Fig. 5; (a) symmetric four-terminal circuit; (b) and (c) asymmetric configurations.
@@ III.2 Three-terminal measurements
First, we discuss a three-terminal measurement in a mirror symmetric geometry, for which we keep the contacts 2, 4 and 6 electrically floating; current flows from contact 1 into contacts 3 and 5. In Fig. 5(a), we present as solid (blue) lines the measured current ratios $-I_{3}(B)/I_{1}(B)$ and $-I_{5}(B)/I_{1}(B)$. As expected, we find that $|I_{3}|$ increases with $B$ while $|I_{5}|$ decreases accordingly, such that $I_{3}(B)+I_{5}(B)=-I_{1}(B)$. The dashed black line indicates the prediction of the model according to Eq. (II.2) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$. For this large mobility, the Ohmic term (first term) of Eq. (II.2) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$), such that with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$ Eq. (4) [Drude model] and Eq. (II.2) [chiral model] become identical.
@@
In the classical limit for $B\lesssim 0.5$ T the Drude model agrees well with our measurement. As we increase $B$ further, the Drude model still predicts the correct trend but as long as we use a constant $\mu(B)=\mu_{0}$ (black dashed lines), we observe growing deviations. These are clearly visible in the enlargement around the first plateau at $\nu=1$ in Fig. 5(b). The red dashed lines are the prediction of the chiral model according to Eq. (4) [or Eq. (II.2) using $\mu B=\infty$], where we inserted a constant filling factor $\nu=1$. It perfectly fits the plateau values of the current ratios. An exact match between measurements and model beyond the plateaus can be achieved by using $\mu(B)$ as a free fit parameter in Eq. (II.2). For the dotted red line in Fig. 5(a), we used the magnetic field dependent values of $\mu(B)$ as shown by red squares in Fig. 6. For comparison, we have also determined $\mu(B)$ from the Shubnikov-de-Hass oscillations of a standard longitudinal resistance measurement, see the black dots in Fig. 6.  [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to $T>4\,$K between the two measurements. The general agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq. (II.2) to our three-terminal current measurements supports the validity of our model.
@@
The two regions of divergent $\mu(B)$ in Fig. 6 correspond to the plateau regions with the local filling factors $\nu=2$ and $\nu=1$. Here, $I_{xx}$ vanishes. Away from the plateau regions, we find strongly reduced mobility values, indicating a transition to diffusive transport. The difference between the Drude model predictions for $\mu=\mu_{0}$ [dashed line in Fig. 5(a)] and the measured currents is the ohmic current contribution $I_{xx}$, which becomes clearly visible for $\mu<20$.
So-far we have discussed three-terminal current measurements. With the knowledge of the contact resistances of our sample, we can quantitatively compare the experiments with model predictions. Between the plateaus, the mobility is reduced and the Drude model fits perfectly to our data, indicating non-chiral transport. Regarding the plateau regions, our three-terminal measurements are clearly consistent with chiral transport. Importantly, any additional non-zero transmission coefficients in Eq. (3) due to non-chiral currents would modify the prediction of Eq. (4). Clearly, we do not observe such deviations.
@@ III.3 Four-terminal measurements
In Fig. 7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements (blue solid lines). Again, we measure the currents into contacts $i=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig. 7(a), contact 6 in (b), and contact 2 in (c). If we assume $\mu(B)=\mu_{0}$, Eq. (II.2) reproduces the overall dependences $I_{i}(B)$ with a similar accuracy and similar deviations as it did for the three-terminal measurement (dashed black lines). The dotted red lines correspond to the predictions of Eq. (II.2), where we used the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf. Fig. 6. As it was the case for the three terminal measurements, for the plateau regions, corresponding to diverging $\mu(B)$, the agreement between theory and the measured currents is perfect. However, away from the plateaus, the theory curves using the varying $\mu(B)$ does no longer yield a perfect agreement between theory and measured data. It means, that the division of the ohmic $I_{xx}$ component of the current is not accurately predicted.
Nevertheless, our four-terminal measurements confirm the main conclusions of our three-terminal measurements: (1) Away from the plateaus, the transport is not chiral. (2) For the plateau regions, the agreement between our four-terminal measurements and the chiral and coherent model supports chiral transport.
III.4 Limits of our model in the diffusive regime
Away from the plateaus, the measured currents depend on the contact configuration beyond the predictions of our model. We conjecture, that our initial assumption of a homogeneous Hall bar fails near the intersections. In the regime of diffusive transport it implies a homogeneous electric field inside the Hall bar. In reality, near intersections, both components, $E_{x}$ and $E_{y}$ of the electric field and, therefore, the currents $I_{xx}$ and $I_{xy}$ become inhomogeneous. While $I_{xy}$ is still completely bent into the contacts, the division of $I_{xx}$ is influenced by its inhomogeneity. A more realistic model would require a suitable numerical calculation, e.g., based on a path integral formalism. We conjecture, that such a corrected model would yield perfect agreement with measurements, namely that the Drude model is still correct for more than three terminals, where it worked perfectly for three terminals.
III.5 Purely experimental confirmation of non-chiral currents between plateaus